An emergency relief airplane is dropping a care package to people on a small island. The plane is flying with a velocity of v0 = 176 km/h at an altitude of h = 393 m.

If the package must land on the island, at what horizontal distance D from the plane to the island should the package be dropped?

Answer:

2.
What is the magnitude of the final velocity of the package the instant before it hits the ground?
Answer #1

use

Time \ t\ taken for an object to fall distance \ d\ : \ t =\ \sqrt {\frac{2d}{g}}

time taken by package to fall 393 meters = 8.95 seconds

initial velocity in x direction of the package = 176 km/h = 48.8889m/s

hence horizontal distance needed = 48.9 * 8.95 = 437 meters

for velocity use

Instantaneous velocity \ v_i\ of a falling object after elapsed time \ t\ :

\ v_i = gt

Vy = 9.8 * 8.95 = 87.71 m/s

in x direction = 48.88 m/s

total velocity = ( x2 + y2)^0.5

=100 m/s